OpenStudy (anonymous):

solve for x: (x-4)^2=2

OpenStudy (bahrom7893):

expand: \[(a-b)^2 = a^2-2ab+b^2\]

OpenStudy (anonymous):

9?

OpenStudy (bahrom7893):

well expand first, what do you get?

OpenStudy (anonymous):

(1-4)^2=1^2-2(1)(4)+4^2..?

OpenStudy (bahrom7893):

\[(x-4)^2 = x^2-8x+16=2\]\[x^2-8x+14=0\]Now just use the discriminant formula.

OpenStudy (anonymous):

whats the discriminant formula??

OpenStudy (bahrom7893):

\[D=b^2-4ac\]\[x_1=\frac{-b-\sqrt{D}}{2a}\]\[x_2=\frac{-b+\sqrt{D}}{2a}\]

OpenStudy (anonymous):

so quadratic pretty much

OpenStudy (bahrom7893):

yea

OpenStudy (anonymous):

\[\large (x-4)^2 = 2\]\[\large \sqrt{(x-4)^2} = \pm \sqrt{2}\]\[\large x-4 = \pm\sqrt{2}\]\[\large x= 4 \pm \sqrt{2}\]

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

thats not one of the options

OpenStudy (anonymous):

you'll have 2 answers. one positive,one negative.

OpenStudy (anonymous):

yeah none of the options look like that though there all written with the positive or negative sign but none of them are that

OpenStudy (anonymous):

Just split it up then....

OpenStudy (anonymous):

4 + ... , 4 - ...

OpenStudy (anonymous):

...oops nevermind it is an option i was looking at a different question..